for real number x and y , defi rne a relationR, xR y if and only if x-y+21/2  is an iirational number.then the relation R is

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find solution to your question below
x, y \in R
Relation is defined
x R y such that
x – y + (2)1/2is an irrational number.
Let a to be an irrational number.
Then,
x - y + \sqrt{2} = a
y = x + \sqrt{2}-a
This is similar to straight line.
So for every input, there is one output
So, R is one – one relation.
Since range is equal to co domain, R is onto also.
Finaaly, R is one – one & onto relation.